# TDE-3: An improved prior for optical flow computation in spiking neural networks

### v1

@@ 3.1 Robust direction-selectivity of individual TDE-3 in textured environments
To compare the direction-selectivity of TDE-2 and TDE-3, we simulated their responses to textures moving in four cardinal directions (Figure 2B): left-to-right (L-R), right-to-left (R-L), top-to-bottom (T-B) and bottom-to-top (B-T). For each stimulus presentation (2000 per epoch), we randomly varied the texture composition, motion direction, and velocity (see Section 5.1 for the details). To show that the robust direction-selectivity of the TDE-3 stems from its structure and is not due to some particular set of parameters, we varied at each epoch (400 in total) all of the TDE parameters over a 10-fold range. Without loss of the generalization, we limited our investigation of direction-selectivity to responses of L-R motion-sensitive TDEs (Figure 2B) that sampled signals from three adjacent pixels through which textures were moving.
Figure 2: Augmented 3-point TDE retains direction-selectivity in a textured environment.A - Stimulus. B. TDE-2 and TDE-3 and their responses to stimuli moving in 4 cardinal directions. C. Direction-selectivity index (fraction of spikes fired upon stimulus motion in PD.
Figure 2B shows how the responses of TDE-2s and TDE-3s depend on the texture motion direction. One can appreciate that while L-R TDE-2 fires to all stimuli directions, the spiking of the TDE-3 is contained to stimuli in its preferred direction. To assess the the TDE’s direction-selectivity quantitatively, we calculated the so-called Direction-Selectivity Index (DSI) as the ratio between number of spikes emitted by motion in PD and the total number of spikes fired by the TDE in an epoch. Figure 2C compares the DSI of TDE-2s and TDE-3s averaged across all of the epochs. For the TDE-2, the mean DSI is 0.36 with a standard deviation of 0.08. This indicates a rather poor direction-selectivity in a textured environment, as on average inference of motion direction with such a detector is only 11% better than random guessing. On the other hand, the TDE-3 has a DSI of exactly 1, with zero deviation. Indicating, that regardless of the initialization of the parameters, simply because of its structure, the TDE-3 is activated only by motion in the PD.
@@ 3.2 Supervised training of TDE-3
To promote linear mapping between the input velocities in PD and the TDE outputs, we performed supervised training with BPTT and surrogate gradients for both spike count (Figure 9) and ISI-based (Figure 10) inference methods and compared obtained results (see Section 2 and 5.2 for the details). Note that in the context of linear mapping of input velocities into TDE outputs, the TDE-2 and TDE-3 are identical because the delay between the activation of the facilitator and trigger is not affected by the inhibitory input. Therefore, although in experiments with synthetic data we perform all of the training on the TDE-3, the results are generalizable to the TDE-2.
There were three types of experiments with training of TDE-3. First, we established that detectors are trainable in principle and that linear mapping between input velocities and TDE output is possible for various stimuli dynamic ranges (Figures 3 and 4, Table 1). Second, we investigated how training and mapping are affected by changes in the spatial frequency of the stimulus - a known liability of correlation-based motion detectors [41] (Figure 5). Third, we studied the robustness of TDE training and inference in the presence of various levels of noise (Figure 6, Table 2)
@@ 3.2.1 Linear mapping of velocity of moving edge
Figure 3: Training of the TDE-3:Wide dynamic range, low resolution. (A)Loss function for spike count-based inference, (B) loss function for ISI-based inference, (C) comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during test; Orange - training with spike count, inference with ISI during test; Green - training with ISI, inference with ISI during test; Red - training with ISI, inference with spike count during test.
@@
Figure 3 shows that the TDE can be trained to encode a wide dynamic range of input velocities. The loss function on Panels A (training with the spike count) and B (training with ISI) gradually decreases during training, indicating that with both inference methods it is indeed possible to train the TDE to discriminate input velocities in a wide dynamic range. Panel C makes a quantitative comparison of velocity coding with various training and inference methods. Unsurprisingly, the correlation coefficient between estimated and true velocity is the highest when evaluation is done with the same inference method as training: r=0.97 for the spike count training and inference (blue line), r=0.99 for the ISI training and inference (green line). Yet training with spike count leads to reasonable performance with ISI inference (r=0.86, orange line) and vice versa (r=0.85, red line) with the only problem being distinguishing between the two highest velocities. Training with ISI-based inference also leads to a lower average number of spikes emitted by stimulus (1.9 vs 4.4), suggesting higher energy efficiency when implemented on neuromorphic hardware [43].
@@
Figure 4 demonstrates that the TDE can effectively encode a narrow dynamic range of input velocities with high precision (Panels A and B). In Panel C, one can appreciate that apart from the case of training with spike count and evaluation with ISI (orange line, $r=0.91$), input velocities are mapped into TDE outputs nearly perfectly, with a correlation $r\approx 1.0$ (Table 1). Also, ISI evaluation with spike count training results in a decrease in dynamic range, as the estimate becomes zero for the first two velocities. This occurs because these two almost equal velocities (0.025 and 0.0256) were encoded with 1 spike, making the ISI indeterminable.
@@
For this experiment, fine resolution was paramount. Therefore, we calculated mean relative error by dividing the difference between true and inferred velocity by true velocity and averaging this metric across all velocities (see Equation 7, Table 1). Unsurprisingly, ISI evaluation with spike count training yielded the largest relative error (30%). Even when excluding the first two velocities, which because TDE emitted only 1 wpike contributed to the variance the most, the relative error remains >= 3 times higher (7.9%) than for the other three conditions, where velocity was encoded with an error as small as 2.1% (ISI training, spike count evaluation, red line). We also note in this experiment that training with ISI led to a higher average number of spikes (7.4) compared to spike count training (5.6) (see Table 1).
@@
Figure 4: Training of the TDE-3: Narrow dynamic range, high resolution.Training of the TDE-3:Wide dynamic range, low resolution. (A)Loss function for spike count-based inference, (B) loss function for ISI-based inference, (C) comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during test; Orange - training with spike count, inference with ISI during test; Green - training with ISI, inference with ISI during test; Red - training with ISI, inference with spike count during test.
@@ 3.2.2 Spatial frequency and velocity coding by TDE
Figure 5: Training of the TDE-3: Robustness to variation in spatial frequency. (A)Loss function for spike count-based inference, (B) Loss function for ISI-based inference. (C)Comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during the test; Orange - training with spike count, inference with ISI during the test; Green - training with ISI, inference with ISI during the test; Red - training with ISI, inference with spike count during the test. As there was no noise, all of the variation in TDE response (shading) originates from the variation in stimulus spatial frequency (spacing between the edges)
@@ 3.2.3 Noise and velocity coding by TDE
Figure 6: Training TDE: Performance in the presence of noise.Top row - low noise, Bottom row - high noise. (A)Loss function for spike count-based inference, (B) loss function for ISI-based inference, (C) comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during test; Orange - training with spike count, inference with ISI during test; Green - training with ISI, inference with ISI during test; Red - training with ISI, inference with spike count during test.
To test whether the training of the TDE to encode velocity is robust towards noise, we injected into our simulation background activity noise - one of the primary sources of noise in event-based vision [16]. The noise was injected as stochastic events with mean rates varied 50-fold in magnitude: from 0.2 Hz/pz to 10Hz/px (see Section 5.1, Table 2). To allow the network to combat noise we employed a Spatio-Temporal Correlation Filter (STCF) of events (Figures 9, 10) that allows an event to pass filtering stage only when there are events in at least n neighboring pixels within a specified timeframe (Section 5.2, [16]) with n being trainable parameter.
@@
Figure 6 shows the results of the training for two extreme cases of "bright" (noise rate=0.2Hz/px) and "dark" (noise rate=5Hz/px) scenes. With the low noise level, the performance of the TDE (Figure 6 top row) was almost identical to the performance in noiseless experiments (Figure 5) in terms of shapes of tuning curves and correlation with stimulus velocities. These observations are also confirmed by FTA (Table 2) with nearly all TDEs activity being caused by the stimulus. Results of the training with a high level of simulated noise are shown in Figure 6, bottom row. With both inference methods, there is an early dip in the loss function (Panels AII and BII) caused by an increase in the number of neighboring pixels where at least one event ought to occur for any given event to pass the STCF processing stage (see Section 5.2). As a result of this, more noise is rejected and loss dramatically decreases.
@@
Panel CII illustrates tuning curves obtained in all 4 testing conditions. One can immediately note that training and inference with ISI (green line) leads to massive standard deviation (shading) in the estimated velocity. Surprisingly, TDEs trained with ISI showed better performance when testing was done with spike count (red line) as the standard deviation was lower and the correlation coefficient was higher (r=0.8 vs r=0.64). With ISI-based training, about half of the encoded motion signals were related to the stimulus (Table 2). Yet, given the low mean spike count (N=1.6) TDEs coding capabilities were rather poor with at least one edge being ignored most of the time. Training with spike count-based inference leads to much better performance both with spike-count (blue line and shading, r=0.83) and ISI-based (orange line and shading, r=0.8) evaluation. Spike count-based training leads to a mean spike count of 7.5 spikes. Since with this training condition almost 70% of encoded motion signals were related to the stimulus (Table 2), we conclude that spike count-based training allows TDE to fairly reliably encode motion signals in the presence of noise.
@@
To summarize, our results indicate that (1) even in the presence of high noise the TDE-3 can be trained to retain its direction-selectivity (Figure 6), (2) the TDE-3 is very robust, with FTA >95% and correlation with stimulus velocity of r $\approx\ 90\%$ for wide range of noise conditions with performance declining only for the two highest noise levels (Table 2), and (3) in the presence of noise the spike-count-based inference serves as a more robust and reliable training signal. The likely reason for this last observation is that ISI dependency on only two spikes makes it susceptible to noise, whereas spike count measures signals over a certain window and therefore mitigates variation in response due to exact spike timing.
Table 2: TDE performance in the presence of noise: summary table.
Noise level Training Inference Test Inference Correlation, r Fraction of True Activity N Spikes
0.2 Hz/px Spike Count Spike Count 0.89 1.0 5.2
ISI 0.86 1.0
ISI Spike Count 0.93 1.0 3.0
ISI 0.97 1.0
0.5 Hz/px Spike Count Spike Count 0.93 0.99 4.8
ISI 0.93 0.99
ISI Spike Count 0.92 0.99 3.0
ISI 0.97 0.99
1.0 Hz/px Spike Count Spike Count 0.91 0.97 5.5
ISI 0.86 0.98
ISI Spike Count 0.92 0.98 3.2
ISI 0.95 0.98
2.0 Hz/px Spike Count Spike Count 0.9 0.96 6.2
ISI 0.91 0.95
ISI Spike Count 0.93 0.96 3.3
ISI 0.96 0.95
5.0 Hz/px Spike Count Spike Count 0.83 0.67 7.5
ISI 0.8 0.63
ISI Spike Count 0.8 0.57 1.6
ISI 0.64 0.52
10.0 Hz/px Spike Count Spike Count 0.61 0.18 1.3
ISI 0.09 0.12
ISI Spike Count 0.66 0.11 5.8
ISI 0.31 0.14
@@ 3.3 Real-world data
To study the performance of the TDE-2 and TDE-3 with real-world data, we used a dataset from [39]. Specifically, we focused on sequence with boxes moving to the right (Figure 7) and disk rotating counter-clockwise (Figure 8). The former is an example of textured stimuli with multiple edges moving in the same direction and thus directly tests the robustness of TDE direction-selectivity and gains obtained by augmenting the TDE-2 with an inhibitory input. The latter contains a high dynamic range of velocities both in magnitude (80-fold) and direction (360°).
The real-world data was used as a test dataset. TDEs were pre-trained on synthetic data (see Sec. 5.2.3) using spike count as the inference method because it tends to give better performance in the presence of noise (Table 2).

### v2

@@ 3.1 Robust direction-selectivity of individual TDE-3 in textured environments
To compare the direction-selectivity of TDE-2 and TDE-3, we simulated their responses to textures moving in four cardinal directions (Figure 2B): left-to-right (L-R), right-to-left (R-L), top-to-bottom (T-B) and bottom-to-top (B-T). For each stimulus presentation (2000 per testing round), we randomly varied the texture composition, motion direction, and velocity (see Section 5.1 for the details). To show that the robust direction-selectivity of the TDE-3 stems from its structure and is not due to some particular set of parameters, we varied at each testing round (400 in total) all of the TDE parameters over a 10-fold range. Without loss of the generalization, we limited our investigation of direction-selectivity to responses of L-R motion-sensitive TDEs (Figure 2B) that sampled signals from three adjacent pixels through which textures were moving.
Figure 2: Augmented 3-point TDE retains direction-selectivity in a textured environment. A - Visual stimulus composed of vertical bars. The bars had three light intensity levels: white, grey, and black. The size of the texture along the motion axis was 80 pixels and 3 pixels along the orthogonal direction. We employed 5 velocities: 0.1 px/timestep, 0.2 px/timestep, 0.33 px/timestep, 0.5 px/timestep and 1 px/timestep. The motion direction (left-right, right-left, top-bottom, bottom-top) and velocity were randomly chosen for each stimulus example (2000 examples per testing round, 400 testing rounds). To vary the "amount" of texture in stimuli we randomly varied the fraction of the grey bars from 0% to 80%. B. TDE-2 and TDE-3 and their responses to stimuli moving in 4 cardinal directions. C. Direction-selectivity index (fraction of spikes fired upon stimulus motion in PD.
Figure 2B shows how the responses of TDE-2s and TDE-3s depend on the texture motion direction. One can appreciate that while L-R TDE-2 fires to all stimuli directions, the spiking of the TDE-3 is contained to stimuli in its preferred direction. To assess the TDE’s direction-selectivity quantitatively, we calculated the so-called Direction-Selectivity Index (DSI) as the ratio between the number of spikes emitted by motion in PD and the total number of spikes fired by the TDE in a testing round. Figure 2C compares the DSI of TDE-2s and TDE-3s averaged across all of the testing rounds. For the TDE-2, the mean DSI is 0.36 with a standard deviation of 0.08. This indicates a rather poor direction-selectivity in a textured environment, as on average inference of motion direction with such a detector is only 11% better than random guessing. On the other hand, the TDE-3 has a DSI of exactly 1, with zero deviation. Indicating, that regardless of the initialization of the parameters, simply because of its structure with inhibitory input that eliminates residual gain activity (details - Section 2, illustration - Figure 1), the TDE-3 is activated only by motion in the PD.
For simplicity of interpretation, in Figure 2 we limited our analysis to 1D motion. Yet, the same observation is true for the 2D motion (Supplementary Figure S1). We performed experiments with textures of increasing complexity (vertical bars, checkerboard, randomly generated textures) moving in 2D and showed that in all of the cases, TDE-3 had higher DSI than TDE-2. However, an increase in texture complexity decreased the DSI of TDE-3 (see Supplementary Figure S1 for the details).
@@ 3.2 Supervised training of TDE-3
To promote linear mapping between the input velocities in PD and the TDE outputs, we performed supervised training with BPTT and surrogate gradients for both spike count and ISI-based inference methods and compared obtained results (see Section 2 and 5.2 for the details). Note that in the context of linear mapping of input velocities into TDE outputs, the TDE-2 and TDE-3 are identical because the delay between the activation of the facilitator and trigger is not affected by the inhibitory input. Therefore, although in experiments with synthetic data we perform all of the training on the TDE-3, the results are generalizable to the TDE-2.
There were three types of experiments with training of TDE-3. First, we established that detectors are trainable in principle and that linear mapping between input velocities in a given cardinal direction and TDE output is possible for various stimuli dynamic ranges (Figures 3 and 4, Table 1). Second, we investigated how training and mapping are affected by changes in the spatial frequency of the stimulus - a known liability of correlation-based motion detectors [50] (Figure 5). Third, we studied the robustness of TDE training and inference in the presence of various levels of noise (Figures 6, 7)
@@ 3.2.1 Linear mapping of velocity of moving edge
Figure 3: Training of the TDE-3:Wide dynamic range, low resolution, one moving edge (A)Loss function for spike count-based inference, (B) loss function for ISI-based inference, (C) comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during the test; Orange - training with spike count, inference with ISI during the test; Green - training with ISI, inference with ISI during the test; Red - training with ISI, inference with spike count during the test.
Figure 3 shows that the TDE can be trained to encode a wide dynamic range of input velocities. The loss function on panels A (training with the spike count) and B (training with ISI) gradually decreases during training, indicating that with both inference methods it is indeed possible to train the TDE to discriminate input velocities in a wide dynamic range.
@@
Panel Figure 3C makes a quantitative comparison of velocity coding with various training and inference methods. Unsurprisingly, the correlation coefficient between estimated and true velocity is the highest when evaluation is done with the same inference method as training: r=0.97 for the spike count training and inference (blue line), r=0.99 for the ISI training and inference (green line). Yet training with spike count leads to reasonable performance with ISI inference (r=0.86, orange line) and vice versa (r=0.85, red line) with the only problem being distinguishing between the two highest velocities. Training with ISI-based inference also leads to a lower average number of spikes emitted by stimulus (1.9 vs 4.4), suggesting higher energy efficiency when implemented on neuromorphic hardware [52].
@@
Figure 4 demonstrates that the TDE can effectively encode a narrow dynamic range of input velocities with high precision (Panels A and B). In Panel C, one can appreciate that apart from the case of training with spike count and evaluation with ISI (orange line, $r=0.91$), input velocities are mapped into TDE outputs nearly perfectly, with a correlation $r\approx 1.0$ (Table 1). Also, ISI evaluation with spike count training results in a decrease in dynamic range, as the estimate becomes zero for the first two velocities. This occurs because these two almost equal velocities (0.0256 and 0.0263) were encoded with 1 spike, making the ISI indeterminable.
@@
For this experiment, fine resolution was paramount. Therefore, we calculated mean relative error by dividing the difference between true and inferred velocity by true velocity and averaging this metric across all velocities (see Equation 7, Table 1). Unsurprisingly, ISI evaluation with spike count training yielded the largest relative error (30%). Even when excluding the first two velocities, which because TDE emitted only 1 spike contributed to the variance the most, the relative error remains >= 3 times higher (7.9%) than for the other three conditions, where velocity was encoded with an error as small as 2.1% (ISI training, spike count evaluation, red line). We also note in this experiment that training with ISI led to a higher average number of spikes (7.4) compared to spike count training (5.6) (see Table 1).
@@
Figure 4: Training of the TDE-3: Narrow dynamic range, high resolution, one moving edge. Training of the TDE-3: wide dynamic range, low resolution. (A)Loss function for spike count-based inference, (B) loss function for ISI-based inference, and (C) comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during the test; Orange - training with spike count, inference with ISI during the test; Green - training with ISI, inference with ISI during the test; Red - training with ISI, inference with spike count during the test.
@@
* Here first two input velocities generated only 1 spike. Therefore the estimated velocity was 0, which led to a large error in the estimated velocity. We excluded those two values from the analysis to get a better fill of what is actual coding error when the velocity can be estimated and get the value of 7.9%
@@ 3.2.2 Spatial frequency and velocity coding by TDE
Figure 5: Training of the TDE-3: Robustness to variation in spatial frequency. Wide dynamic range, low resolution, two moving edges at randomly assigned distances. (A)Loss function for spike count-based inference, (B) Loss function for ISI-based inference. (C)Comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during the test; Orange - training with spike count, inference with ISI during the test; Green - training with ISI, inference with ISI during the test; Red - training with ISI, inference with spike count during the test. As there was no noise, all of the variation in TDE response (shading) originates from the variation in stimulus spatial frequency (spacing between the edges)
@@ 3.2.3 Noise and velocity coding by TDE
Figure 6: Training TDE: Performance in the presence of noise. Top row - low noise, Bottom row - high noise. (A)Loss function for spike count-based inference, (B) loss function for ISI-based inference, and (C) comparison of the velocity tuning curves. Blue - training with spike count, inference with spike count during the test; Orange - training with spike count, inference with ISI during the test; Green - training with ISI, inference with ISI during the test; Red - training with ISI, inference with spike count during the test.
To test whether the training of the TDE to encode velocity is robust towards noise, we injected into our simulation background activity noise - one of the primary sources of noise in event-based vision [18]. The noise was injected as stochastic events with mean rates varied 50-fold in magnitude: from 0.2 Hz/px to 10Hz/px (see Section 5.1).
To allow the network to combat noise we employed a Spatio-Temporal Correlation Filter (STCF) of events (Section 5). STCF is a popular noise-combating tool in the field of event-based vision [18, 48, 2] that allows an event to pass the filtering stage only when there are events in at least n neighboring pixels within a specified timeframe (Section 5.2, [18]) with n being a trainable parameter. Although as a result we would, strictly speaking, test not the performance of the TDE, but of the TDE and STCF, our approach is still sound. The reason for this is that STCF is almost always part of the TDE circuitry anyhow [29, 30, 32, 33, 35]. Therefore, since in real applications TDE will be always paired with STCF, it makes sense to investigate their joint performance. Especially, given that STCF does not eliminate all of the noise and there is always a concern that aggressive filtering would destroy part of the signal.
@@
Figure 6 shows the results of the training for two extreme cases of "bright" (noise rate=0.2Hz/px) and "dark" (noise rate=5Hz/px) scenes. With the low noise level, the performance of the TDE (Figure 6 top row) was almost identical to the performance in noiseless experiments (Figure 5) in terms of shapes of tuning curves and correlation with stimulus velocities. These observations are also confirmed by FTA (Figure 7) with nearly all TDEs activity being caused by the stimulus. Results of the training with a high level of simulated noise are shown in Figure 6, bottom row. With both inference methods, there is an early dip in the loss function (Panels AII and BII) caused by an increase in the number of neighboring pixels where at least one event ought to occur for any given event to pass the STCF processing stage (see Section 5.2). As a result of this, more noise is rejected and loss dramatically decreases.
@@
Panel CII illustrates tuning curves obtained in all 4 testing conditions. One can immediately note that training and inference with ISI (green line) leads to massive standard deviation (shading) in the estimated velocity. Surprisingly, TDEs trained with ISI showed better performance when testing was done with spike count (red line) as the standard deviation was lower and the correlation coefficient was higher (r=0.8 vs r=0.64). With ISI-based training, about half of the encoded motion signals were related to the stimulus (Figure 7). Yet, given the low mean spike count (N=1.6), TDEs coding capabilities were rather poor with at least one edge being ignored most of the time. Training with spike count-based inference leads to much better performance both with spike-count (blue line and shading, r=0.83) and ISI-based (orange line and shading, r=0.8) evaluation. Spike count-based training leads to a mean spike count of 7.5 spikes. Since with this training condition almost 70% of encoded motion signals were related to the stimulus (Figure 7), we conclude that spike count-based training allows TDE to fairly reliably encode motion signals in the presence of noise.
@@
In Figure 7 we summarized the influence of the noise levels on the correlation between output and ground truth input velocity (panel A), FTA (panel B), and mean number of spikes generated by a moving edge (panel C). Panels A and B show that TDE-3 is very robust up to a noise level of up to 2Hz/px with FTA >95% and correlation with the stimulus velocity r of $\approx\ 90\%$ throughout and largely regardless of the training / inference method. However, a further increase in noise level led to a sharp decrease in performance with most of the TDE output being noise-driven. Our results also show that in the presence of noise, the spike-count-based inference serves as a more robust and reliable training signal. The likely reason for this is that ISI dependency on only two spikes makes it susceptible to noise, whereas spike count measures signals over a certain time window, and therefore mitigates variation in response caused by exact spike timing.
Figure 7: Training of the TDE-3: Summary statistics of TDE performance in the presence of noise. A: Correlation of the TDE output with ground truth input velocity as a function of noise magnitude. B: FTA as a function of noise magnitude. C: Average number of spikes used for encoding of an edge velocity as a function of noise level. In all three cases, noise levels are plotted on the logarithmic scales. For panels A and B the color code is as follows: Blue - training with spike count, inference with spike count during the test; Orange - training with spike count, inference with ISI during the test; Green - training with ISI, inference with ISI during the test; Red - training with ISI, inference with spike count during the test. For panel C blue describes training with spike count, green - training with ISI.
@@ 3.3 Real-world data
To study the performance of the TDE-2 and TDE-3 with real-world data, we used a dataset from [48]. Specifically, we focused on sequence with boxes moving to the right (Figure 10) and disk rotating counter-clockwise (Figure 11). The former is an example of textured stimuli with multiple edges moving in the same direction and thus directly tests the robustness of TDE direction-selectivity and gains obtained by augmenting the TDE-2 with an inhibitory input. The latter contains a high dynamic range of velocities both in magnitude (80-fold) and direction (360°) and is used to probe velocity coding by TDEs in real-world scenes.
The real-world data was used as a test dataset. TDEs were pre-trained on synthetic data (see Sec. 5.2.3) using spike count as the inference method because it gives better performance in the presence of noise (Figure 7). We pooled together ON and OFF events as it is often encountered in biological systems [15, 54, 55].